A car starts from stoplight and is traveling with a velocity of 10m/second east in 20 seconds. What is the acceleration of the car?


1
Expert's answer
2021-06-28T23:37:02-0400

Solution.

"v_0=0;"

"v=10m\/s;"

"t=20s;"

"a=\dfrac{v-v_0}{t};"

"a=\dfrac{10m/s-0}{20s}=0.5v/s^2;"

Answer: "a=0.5m\/s^2."